1.Edmund buys a television set from a merchant who ask $1,250.00 at the end of 60 days. Edmund wishes to pay immediately and the merchant offers to compute the cash price on the assumption that money is worth 8% simple interest. What is the cash price?
i really need your help..pleeassssssssssse.thanks

Answer by Maths68(1474)

Let
Cash price = x
8% interest on Cash price = x * 8% = x*8/100 = x * 2/25 = 2x/25
Total cost including interest = 1250
Cash price + 8% interest on Cash price = Total cost including interest
x+(2x/25)=1250
Multiply by 25 both sides of the above equation
25x+2x=31250
27x=31250
Divide by 27 both sides of the above equation
27x/27=31250/27
x=1157.41
Cash price = $1157.41